given the utility function U=150X+40X2-X3, derive average and marginal utility functions, Find the value of X at which total utility is maximum, and the value of X at which average utility is maximum?


1
Expert's answer
2021-11-22T10:02:34-0500

Average utility


"\frac{dU}{dX} = 150 + 80x -3x^2"


"3X^2 -80X -150 = 0"


"x(3X- 4 0) = 0, ( 40X -150) = 0"


Value of X when average utility is maximum

"3x^2 - 40X =0"


"3X^2 = 40X"


"3X = 40"


"x =\frac {40}{3} = 13.33 maximum"


Also

"40X -150 = 0"


"40X = 150"


"x = 3.75 minimum"



Marginal utility


"\frac{dU}{dX} = 150 + 80x -3x^2"


"\frac {dU}{dX} = 80-6X = 0"


"X =\frac {80}{6} = 13.3"


Value of X when utility is maximum


"X = 13.3"